# Cross-Modal Temporal Fusion for Financial Market Forecasting

### v1

@@ 5.1. Dataset
To train CMFT, the data is partitioned chronologically into distinct splits: 804 training days (02/2019 – 07/2022), 268 validation days (08/2022 – 09/2023), and 268 test days (10/2023 – 05/2024). The final tensor structure preserves cross-modal interactions between market movements (price), macro-indexes (bond/GDP/CPI), and corporate disclosures (news/reports).
5.2. Configuration
All experiments were performed on an Nvidia GeForce RTX 4060 laptop GPU with CUDA version 12.6. The lookback window $b$ for the weighted moving average $WMA_{t}$ is set to 30, while the temporal feature expansion window $T^{{}^{\prime}}$ is 90. The five-dimensional rating vector comprises Risk, Market Conditions, Regulation, ESG, and Innovation, with rating scores ranging from 1 to 9, extracted from Llama3. Missing values are handled using linear interpolation for numerical data, while zero-imputation is applied to text embeddings.
To optimize the Transformer model, we employ Optuna (Akiba et al., 2019) for efficient hyperparameter tuning, covering both the model architecture and the training parameters.
For the model architecture, the embedding dimension is selected from $\{32,64,128,256,512,1024\}$. The number of attention heads is chosen from $\{2,4,8,16\}$, ensuring divisibility by $d_{\text{model}}$. The model consists of multiple transformer encoder layers, with num_layers set from $\{1,2,4,8\}$. The FFN layer dimension is optimized from $\{256,512,1024,2048,4096\}$.
[…]
For the baselines, ARIMA was configured per target with automatic order selection. LSTM used sequential inputs with 50 units, ReLU activation, and 200 training epochs. SVR is using a linear kernel and trained separate models for each target variable, with an average result calculated.
Table 3. Performance comparison for classification across baselines and CMTF.
@@ 5.2. Configuration
5.3. Baseline Comparison Models
@@
Linear Regression: Models linear relationships between dependent and independent variables by minimizing the sum of squared residuals to fit an optimal hyperplane (Weisberg, 2005).
ARIMA: Combines autoregressive (AR), differencing (I), and moving average (MA) components to capture temporal dependencies, trends, and seasonality in non-stationary time series (Box et al., 2015).
@@
Random Forest: An ensemble method that aggregates predictions from multiple decorrelated decision trees, reducing overfitting via bootstrap aggregation and feature randomization (Breiman, 2001).
Support Vector Regression (SVR): Extends support vector machines to regression tasks by mapping inputs to a high-dimensional space and optimizing a margin-sensitive loss function (Drucker et al., 1996).
@@
LSTM: A recurrent neural network variant with gating mechanisms (input, output, forget gates) to model long-term sequential dependencies while mitigating vanishing gradients (Hochreiter and Schmidhuber, 1997).
Encoder-only transformer: Adapts self-attention (Vaswani et al., 2017) mechanisms for time series by encoding positional information and temporal relationships, following techniques in (Devlin et al., 2019).
5.4. Evaluation Metrics
Following the approaches taken in previous studies (Shobayo et al., 2024; Singh and Khushi, 2021; Pei et al., 2025; Deng et al., 2019; Sawhney et al., 2020), we assess our result using Precision, Recall, and F1 score to evaluate model performance. For all three metrics, higher values indicate better model performance.
(22) $\text{Precision}=\frac{\text{TP}}{\text{TP}+\text{FP}}$
(23) $\text{Recall}=\frac{\text{TP}}{\text{TP}+\text{FN}}$
[…]
In our classification scheme: true positive (TP) indicates we correctly predicted an increase in price; true negative (TN) indicates we correctly predicted a decrease in price; false positive (FP) indicates we incorrectly predicted a price increase when the price decreased; and false negative (FN) indicates we incorrectly predicted a price decrease when the price increased.
Note that we focus only on classification metrics and deliberately avoid using error metrics that are often applied for regression tasks. This is because a zero change model ($p_{t+1}=p_{t}$) often returns a low root mean squared error (RMSE) or mean absolute percentage error (MAPE). Appendix A shows that the zero change model has the lowest RMSE and MAPE of all models analysed.
Table 4. Ablation study on CMTF with different configurations (+/-, I/N/R), representing with/without tensor interpreting (I), news data (N), and financial report data (R).
@@ 5.4. Evaluation Metrics
5.5. RQ1 Performance Comparison
Table 3 presents a comprehensive performance evaluation of our proposed CMTF against different baselines. To test the effectiveness of our framework, we do not enable the Tensor Representation module: we only include Tensor Encoding, Tensor Interpreting, and Transformer forecasting in the classification settings.
In classification, our proposed CMTF framework exhibits the highest recall of 84.88% and an F1-score of 0.64, outperforming all baselines. This highlights its strength in capturing sequential dependencies and leveraging multimodal data sources to improve predictive accuracy. Random Forest achieves an F1-score of 0.60, indicating its effectiveness in feature selection and ensemble learning, though its recall 71.10% remains lower than CMTF.
5.6. RQ2 Ablation Study
5.7. RQ3 Module Sensitivity
Table 5. Ablation study comparing base methods with Tensor Interpreting (+ I).
@@ 5.7. RQ3 Module Sensitivity
To better understand the impact of Tensor Interpreting (+I), we conduct an ablation study to compare various baseline methods with and without this component. From Table 5, we find that the impact of +I is nuanced. Precision remains relatively stable across models, indicating that Tensor Interpreting does not compromise class separability. Instead, recall exhibits noticeable improvements, particularly for Transformer and LSTM, suggesting that +I aids in identifying more relevant patterns for classification.
The results confirm that Tensor Interpretation enhances classification performance, particularly for models that rely heavily on feature transformations (e.g., SVR, LSTM, Transformer).
6. Discussion
Although we have evaluated our framework across multiple time series forecasting models, several challenges remain. One key limitation is the absence of a publicly available standardized dataset for multimodal and cross-modal financial forecasting. For example, in (Mandal et al., 2024), they used data from (Xu and Cohen, 2018), which includes only tweets and historical prices, limiting its applicability to broader multimodal scenarios. Although (Lim et al., 2021) discussed methods for handling different data granularities, it does not address the integration of unstructured data, such as textual information, into forecasting models. Another major challenge is data privacy. In a previous study (Cheng et al., 2022), they compiled a diverse dataset that included financial events, news, historical prices, and knowledge graph data. However, due to privacy concerns, the dataset was not made publicly available, restricting reproducibility and benchmarking opportunities for future research.
@@ 6. Discussion
7. Conclusion
We introduce Cross-Modal Temporal Fusion (CMTF), a transformer-based deep learning framework for financial market forecasting. To effectively capture the interactions between historical price trends, macro indexes, and textual financial data, CMTF incorporates specialized components. Specifically, these include: (1) an attention-based cross-modal fusion mechanism that dynamically weighs the contribution of different modalities, (2) a tensor interpretation module to extract relevant cross-modal features, and (3) an auto-training scheme to streamline model iteration and optimization. Using real-world financial data sets, we demonstrate that CMTF outperforms all baselines on the classification task. Lastly, we explore the interpretability of our model, highlighting how CMTF can (i) analyze the relative importance of different modality data and (ii) adapt to evolving market dynamics through its feature selection mechanisms. For industrial users, CMTF is more than just a financial market forecasting model; it serves as a robust framework to efficiently handle multimodal data.
Acknowledgements
@@
Appendix A Appendix: Regression Analysis
Here we provide regression analysis of the models. The regression task focuses on predicting the next day’s closing price for all $N$ stocks. Given the input tensor $\mathcal{X}\in\mathbb{R}^{T\times D}$, the model aims to predict the future stock price at time $t+1$. The objective is to minimize the mean squared error between the predicted and actual stock prices:
(25) $\hat{P}_{t+1}=f(\mathcal{X}_{T})$
where $\hat{P}_{t+1}$ is a vector of predicted multi-stock closing prices at time $(t+1)$, and $f$ represents the learned function. The goal is to accurately forecast the next closing price of all $N$ stocks.
We use the widely adopted error-based metrics RMSE and MAPE, where lower values indicate better predictive accuracy. RMSE and MAPE are defined as:
(26) $\text{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(\hat{p}^{i}_{t+1}-{p}^{i}_{t+1})^{2}}$
(27) $\text{MAPE}=\frac{1}{n}\sum_{i=1}^{n}\left|\frac{\hat{p}^{i}_{t+1}-{p}^{i}_{t+1}}{{p}^{i}_{t+1}}\right|\times 100$
Results presented in Table 6 show that the Zero Change model has the lowest RMSE and MAPE. Despite this, the Zero Change model has very poor classification accuracy, with precision and recall both less than 49% (see Table 3). This demonstrates that RMSE and MAPE are unsuitable evaluation metrics for models predicting tomorrow’s price movement.
Table 6. Performance comparison for regression task across baselines and CMTF.
Zero Linear ARIMA RF SVR LSTM CMTF
RMSE $\downarrow$ 1.07 1.31 5.54 2.79 1.21 9.25 3.35
MAPE (%) $\downarrow$ 2.05 3.72 20.91 9.08 3.53 33.78 10.04

### v2

@@ 5.1 Dataset
To train CMTF, the data is partitioned chronologically into distinct splits: 804 training days (02/2019 – 07/2022), 268 validation days (08/2022 – 09/2023), and 268 test days (10/2023 – 05/2024). The final tensor structure preserves cross-modal interactions between market movements (price), macro-indexes (bond/GDP/CPI), and corporate disclosures (news/reports). For more details, please refer to our GitHub.
5.2 Configuration
All experiments were performed on an Nvidia GeForce RTX 4060 laptop GPU with CUDA version 12.6. The lookback window $b$ for the weighted moving average $WMA_{t}$ is set to 30, while the temporal feature expansion window $T^{{}^{\prime}}$ is 90. The five-dimensional rating vector comprises Risk, Market Conditions, Regulation, ESG, and Innovation, with rating scores ranging from 1 to 9, extracted from the pretrained LLM (Llama-3.1-8B). Missing values are handled using linear interpolation for numerical data, while zero-imputation is applied to text embeddings.
To optimize the Transformer model, we employ Optuna [1] for efficient hyperparameter tuning, covering both the model architecture and the training parameters. For the model architecture, the embedding dimension is selected from $\{32,64,128,256,512,1024\}$. The number of attention heads is chosen from $\{2,4,8,16\}$, ensuring divisibility by $d_{\text{model}}$. The model consists of multiple transformer encoder layers, with num_layers set from $\{1,2,4,8\}$. The FFN layer dimension is optimized from $\{256,512,1024,2048,4096\}$.
For training, the learning rate is searched within $\{$1e-5, 5e-5, 1e-4, 5e-4, 1e-3, 5e-3, 1e-2}\}. The batch size is chosen from $\{32,64,128\}$, balancing computational efficiency and model convergence. The number of training epochs is set from $\{10,20,50,100\}$ to regulate training duration and stability. For baselines, ARIMA was configured with automatic order selection, which finds the best combination of ($p$, $d$, $q$) by evaluating multiple models; LSTM used sequential inputs with 50 units, ReLU activation, and 200 training epochs; and SVR used a linear kernel with separate models trained for each target variable, with an average result calculated.
Table 3: Classification performance comparison.
@@ 5.2 Configuration
5.3 Baseline Comparison Models
@@
Linear Regression: Models linear relationships between dependent and independent variables by minimizing the sum of squared residuals to fit an optimal hyperplane [38].
ARIMA: Combines autoregressive (AR), differencing (I), and moving average (MA) components to capture temporal dependencies, trends, and seasonality in non-stationary time series [3].
@@
Random Forest: An ensemble method that aggregates predictions from multiple decorrelated decision trees, reducing overfitting via bootstrap aggregation and feature randomization [4].
Support Vector Regression (SVR): Extends support vector machines to regression tasks by mapping inputs to a high-dimensional space and optimizing a margin-sensitive loss function [11].
@@
LSTM: A recurrent neural network variant with gating mechanisms (input, output, forget gates) to model long-term sequential dependencies while mitigating vanishing gradients [14].
Encoder-only transformer: Adapts self-attention [35] mechanisms for time series by encoding positional information and temporal relationships, following techniques in [10].
5.4 Evaluation Metrics
Following the approaches taken in previous studies [29, 31, 23, 9, 28], we assess our result using Precision, Recall, and F1 score to evaluate model performance. For all three metrics, higher values indicate better model performance.
$\text{Precision}=\frac{\text{TP}}{\text{TP}+\text{FP}}$ (20)
$\text{Recall}=\frac{\text{TP}}{\text{TP}+\text{FN}}$ (21)
$\text{F1 Score}=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$ (22)
In our classification scheme: true positive (TP) indicates we correctly predicted an increase in price; true negative (TN) indicates we correctly predicted a decrease in price; false positive (FP) indicates we incorrectly predicted a price increase when the price decreased; and false negative (FN) indicates we incorrectly predicted a price decrease when the price increased. Note that we focus only on classification metrics and deliberately avoid using error metrics that are often applied for regression tasks. This is because a zero change model ($p_{t+1}=p_{t}$) will return a low root mean squared error or mean absolute percentage error.
Table 4: Ablation study on CMTF with different configurations (+/-, I/N/R), denoting with/without tensor interpreting (I), news (N), and financial reports (R).
@@ 5.4 Evaluation Metrics
Table 5: Ablation study comparing base methods with Tensor Interpreting (+ I).
@@
5.5 RQ1 Performance Comparison
Table 3 presents a comprehensive performance evaluation of our proposed CMTF against different baselines. To test the effectiveness of our framework, we do not enable the Tensor Representation module: we only include Tensor Encoding, Tensor Interpreting, and Transformer forecasting in the classification settings. In classification, our proposed CMTF framework exhibits the highest recall of 84.88% and an F1-score of 0.64, outperforming all baselines. This highlights its strength in capturing sequential dependencies and leveraging multimodal data sources to improve predictive accuracy. Random Forest achieves an F1-score of 0.60, indicating its effectiveness in feature selection and ensemble learning, though its recall 71.10% remains lower than CMTF.
5.6 RQ2 Ablation Study
5.7 RQ3 Module Sensitivity
To understand the impact of Tensor Interpreting (+I), we conduct an ablation study to compare baseline methods with and without this component. From Table 5, we find that the impact of +I is nuanced. Precision remains relatively stable across models, indicating that Tensor Interpreting does not compromise class separability. In contract, recall exhibits noticeable improvements, particularly for Transformer and LSTM, suggesting that +I aids in identifying more relevant patterns for classification. The results confirm that Tensor Interpretation enhances performance, particularly for models that rely heavily on feature transformations (e.g., SVR, LSTM, Transformer).
6 Discussion
Although we have evaluated our framework across multiple time series forecasting models, several challenges remain. One key limitation is the absence of a publicly available standardized dataset for multimodal and cross-modal financial forecasting. For example, in [21], they used data from [41], which includes only tweets and historical prices, limiting its applicability to broader multimodal scenarios. Although [19] discussed methods for handling different data granularities, it does not address the integration of unstructured data, such as textual information, into forecasting models. Another major challenge is data privacy. In a previous study [5], they compiled a diverse dataset that included financial events, news, historical prices, and knowledge graph data. However, due to privacy concerns, the dataset was not made publicly available, restricting reproducibility and benchmarking opportunities for future research.
@@ 6 Discussion
7 Conclusion
We introduce Cross-Modal Temporal Fusion (CMTF), a transformer-based deep learning framework for financial market forecasting. To effectively capture the interactions between historical price trends, macro indexes, and textual financial data, CMTF incorporates specialized components. These include: (1) an attention-based cross-modal fusion mechanism that dynamically weighs the contribution of different modalities, (2) a tensor interpretation module to extract relevant cross-modal features, and (3) an auto-training scheme to streamline model iteration and optimization. Using real-world financial data sets, we demonstrate that CMTF outperforms all baselines on the classification task. Lastly, we explore the interpretability of our model, highlighting how CMTF can (i) analyze the relative importance of different modality data and (ii) adapt to evolving market dynamics through its feature selection mechanisms. For industrial users, CMTF is more than just a financial market forecasting model; it serves as a robust framework to efficiently handle multimodal data.
{ack}